### v1

# Electrical Control of Optically Active Single Spin Qubits in ZnSe

Abstract
Electrons bound to shallow donors in ZnSe quantum wells are promising candidates for optically addressable spin qubits and single-photon sources. However, their optical coherence and indistinguishability are often limited by spectral broadening arising from charge fluctuations in the local environment. Here, we report electrical control of single donor qubits in ZnSe quantum wells. The applied field induces a DC Stark shift that tunes the emission energy over a range exceeding 30 times the inhomogeneous linewidth, effectively compensating for emitter-to-emitter variations. Concurrently, the field stabilizes trap occupancy, yielding a twofold reduction in optical linewidth and the suppression of spectral wandering. A statistical model based on trap dynamics qualitatively reproduces these observations and elucidates the mechanism of field-assisted charge noise suppression. Our results identify electrical control as a versatile pathway to significantly improve optical and spin addressability.
I Introduction
Optically addressable qubits in semiconductors are attractive building blocks for quantum network and quantum memory applications [doi:10.1126/science.abg1919, Knaut2024Entanglement, awschalom2025challenges]. Among them, the coupled electron-nuclear spin system associated with the donors offers a multilevel platform with potentially long coherence times [Linpeng2018Coherence, Karin2016Longitudinal, Viitaniemi2022Coherent]. Direct bandgap host crystals such as ZnSe provide efficient radiative transitions [Qiao2024, Kutovyi2022Efficient, Kim2012Semiconductor] and, through isotopic purification [Pawlis2019MBE, Kopteva2019-yp, Kirstein2021Extended], an almost nuclear spin free environment for the donors. In particular, shallow donors in ZnSe quantum wells have emerged as a promising platform [De-Greve2010-fi, Karasahin2022, doi:10.1021/nl303663n]. However, their potential is limited due to the environmental decoherence [Kuhlmann2013ChargeSpinNoise, Kim2014-od]. Static disorder within the host material gives rise to inhomogeneous broadening, while dynamic charge noise, originating from charge carriers trapped at nearby defects and interfaces, induces spectral diffusion via the Stark effect [Jiang2024]. These processes introduce dephasing channels that degrade photon indistinguishability, thereby limiting Hong-Ou-Mandel visibility [Sanaka2012Entangling, Sanaka2009Indistinguishable] and reducing spin coherence times.
As a solution, electrical control provides a versatile approach to tune the optical properties of semiconductor quantum emitters [doi:10.1126/science.278.5346.2114]. In particular, applied electric fields can mitigate inhomogeneous broadening caused by variations in local strain and charge environment by tuning disparate emitters into resonance via the DC Stark effect [Patel2010RemoteHOM, PhysRevLett.108.043604, bushmakin2025twophotoninterferencephotonsremote, PhysRevLett.131.033606, zhai2022quantum]. Furthermore, external electric fields can stabilize the fluctuations in local electric field, reducing spectral wandering caused by charge noise [Anderson2019Electrical, Somaschi2016NearOptimal]. These fields are commonly applied using device structures such as p-i-n diodes [Kuhlmann2015TransformLimited, PhysRevLett.104.047402, zeledon2025minute], Schottky diodes [steidl2025single], or surface electrodes [Aghaeimeibodi2021Electrical, DeSantis2021Stark]. Despite these advantages, the difficulty of p-doping ZnSe has prevented the realization of electric-field control for single impurities [PhysRevLett.74.1131, PhysRevB.45.10965, doi:10.1021/acsaelm.4c01104].
In this Letter, we demonstrate electrical control of single donor-bound excitons in ZnSe quantum wells using laterally applied fields from patterned surface electrodes. This configuration enables local tuning of the emission energy by polarizing the bound exciton. In addition to spectral tuning, we find the lateral field actively stabilizes the local charge environment, manifesting as a two fold narrowing of the emission linewidth. We attribute this narrowing to the suppression of spectral wandering arising from trap-induced charge fluctuations, a mechanism corroborated by similar behavior under a weak, above-band optical pump. A statistical model based on trap charge dynamics qualitatively reproduces the observed line broadening and subsequent narrowing under both optical and electrical stabilization. These results establish electric fields as a powerful tool for simultaneously achieving precise spectral tuning and stabilizing the charge environment of quantum emitters.
II Initial Characterization
Fig. 1(a) provides an illustration of our device. The initial sample consists of a GaAs‐substrate–based ZnMgSe/ZnSe/ZnMgSe quantum well incorporating a Cl $\mathrm{\delta}$-doped layer at its center [Ohkawa1987]. Based on first-principles calculations and growth conditions, we expect the majority of chlorine atoms to occupy substitutional Se sites, acting as shallow donors $Cl_{Se}^{+1}$ in the ZnSe lattice, as illustrated in Fig. 1(b) [dosSantos2011, Poykko1998, Wu2022, Wu2022NativeZnSe, 10.1063/5.0265591]. The donor-bound electron acts as a spin $1/2$ qubit and is optically accessible through radiative recombination of bound exciton [Dean1981-sh, Steiner1985-ti, Merz1972OpticalProperties]. To apply a lateral electric field across individual impurity sites distributed inside well, we pattern interdigitated electrodes via a standard metal liftoff process (see the Methods section A) on the surface. A potential difference is applied at two opposing electrodes. The top $\mathrm{AlO}_{\mathrm{x}}$ layer prevents current flow through the quantum well, suppressing leakage current and Joule heating. Fig. 1(c) shows an optical microscope image of the device from the top.
Figure 1: Device illustration and magneto spectroscopy characterization. (a) Schematic of device geometry, consisting of ZnMgSe/ZnSe:Cl/ZnMgSe quantum well, ZnSe buffer, AlOx capping and GaAs substrate, with electrodes on the top and connections. (b) Crystal structure Zn(blue)Se(pink) with Cl(green) as a substitutional donor and bound electron (yellow). (c) Optical microscope image of the fabricated electrodes. The gold-colored regions highlight the patterned electrodes, designed for lateral electric field application. (d) Photoluminescence spectrum of donor-bound exciton in Voigt configuration shows four distinct peaks with two orthogonal polarization.
[…]
We initially characterize the device using photoluminescence spectroscopy under above-band excitation at $405\text{\,}\mathrm{nm}$. We scan the regions between the electrodes to locate photoluminescent impurity sites. At various sites, we observe isolated sharp emission peaks corresponding to $D^{0}X$ to $D^{0}$ transition of single Cl donors. Previous studies on the same quantum emitter have shown antibunching from this transition [Karasahin2022, Jiang2024, Kutovyi2022Efficient].
To confirm the origin of the emission, we perform polarization-resolved magnetospectroscopy. Fig. 1(d) shows the photoluminescence signal from the emitter for magnetic fields up to $\T9\text{\,}\T$ perpendicular to the sample growth direction (Voigt configuration). The external field lifts the degeneracy of the electron and hole spin states in the bound exciton complex and produces four allowed optical transitions [astakhov2002binding]. Each pair of transitions has a specific linear polarization, and together they form two orthogonal polarization bases, shown as H and V [Greilich2012-ic, De-Greve2010-fi]. From the Zeeman-split lines we extract effective g-factors corresponding to each polarization (see Supplemental Material section I). Based on the anticipated level structure, we find the ground state electron g-factor of $g_{e}=1.19\pm 0.02$, and excited state heavy hole g-factor of $g_{h}=0.09\pm 0.02$. Both values are in close agreement with previous studies on single Fluorine donors [Greilich2012-ic, Kopteva2019-yp, De-Greve2010-fi]. For all subsequent measurements, we focus on the electrical properties of this transition at zero magnetic field.
III Spectral Tuning
Fig. 2(a) shows the response of the bound exciton line as a function of applied bias voltage. The emission undergoes a voltage-dependent spectral shift towards longer wavelengths. We attribute this shift to the quantum-confined Stark effect [fu1987electric], in which the applied electric field polarizes the bound exciton by spatially separating the electron and hole [brum1985electric], inducing a dipole moment and reducing the recombination energy, resulting in a redshift of the emission [miller1985electric, PhysRevLett.53.2173]. The dashed black line is a numerical fit to the central wavelength of the emission signal, containing polynomial terms up to fourth degree. We performed this fit by initially fitting a Voigt profile to each spectrum, extracting emission properties including central wavelength, total intensity and linewidth. From this fit, we obtained a polarizability of $\beta=(2.6\pm 0.2)\times 10^{-6}\ \mathrm{meV}\,(\mathrm{cm}/\mathrm{kV})^{2}$ and a permanent dipole moment of $d=(1.0\pm 0.8)\times 10^{-4}\ \mathrm{meV}\,(\mathrm{cm}/\mathrm{kV})$. The first order contribution is relatively small, showing the lack of a notable permanent dipole moment. Additionally, we find $(4.1\pm 1.1)\times 10^{-10}\ \mathrm{meV}\,(\mathrm{cm}/\mathrm{kV})^{3}$ and $(1.1\pm 0.2)\times 10^{-12}\ \mathrm{meV}\,(\mathrm{cm}/\mathrm{kV})^{4}$ for higher-order dependency of Stark shift as a function of the electric field. From this analysis, we can deduce that the quadratic and quartic terms are dominant, which we attribute to the centrosymmetric nature of excitonic transitions in a quantum well [mendez1989optical]. Details of our finite-element simulations used to calculate the internal field from applied voltage bias are provided in the Supplemental Material section IV.A.
Figure 2: Characterization of the Stark shift in a single bound exciton. (a) Photoluminescence bias map. (b) Integrated intensity as a function of the applied bias. (c) Full width at half maximum of the emission line as a function of the applied bias.
Figure 3: Manipulation of charge environment around donor site. (a) Emission spectrum under resonant excitation of the free exciton line for different electric fields. (b) The narrowing effect from the electric field suggests an optimal field of around 35 V is suitable to stabilize the bound exciton emission. The theoretical curve corresponds to about 18 traps distributed in the region of $3\text{\,}\mathrm{n}\mathrm{m}$ to $5\text{\,}\mathrm{n}\mathrm{m}$ from donor with average trap depth of $0.3\text{\,}\mathrm{e}\mathrm{V}$ and initial occupancy of 0.35. We extract these values based on a numerical fit constrained by the donor’s effective Bohr radius and typical defect densities in ZnSe. (c) Emission spectra recorded for varying weak above-band excitation powers in the absence of an electric field. (d) Dynamics of linewidth and central wavelength as a function of different above-band intensities. The theoretical curves correspond to an average saturation power of 1 nW and an initial trap occupancy of about 0.4. We extract these values based on a numerical fit to a statistical model. (e) The electric field stabilizes the fluctuating traps by polarizing the occupancy probability toward 0. (f) The process of filling trap states and eliminating the fluctuating charge environment through free carrier injection pushes occupancy toward unity.
We next examine how the electric field affects the brightness of the impurity-bound exciton. Fig. 2(b) plots the total intensity of emission as a function of voltage bias. For the absolute biases lower than 40 V, the intensity level is stable, indicating that the electric field has a negligible destructive effect on the emission. At higher fields, we observe a decrease in the intensity, indicating a quenching behavior which determines the fundamental tuning range of our device. We attribute the quenching to tunneling of the charge from the impurity site and dissociation of the bound exciton in high fields [PhysRevLett.55.2610, wu1988exciton]. Supplemental Material section IV.B discusses the dynamics of free-exciton and free-trion emission under an electric field, showing that the bound exciton is significantly more stable than the negatively charged trion due to the strong confinement of the ground-state electron at the impurity site [PhysRevB.55.R1970].
In the next step, we investigate the behavior of the linewidth as a function of the applied electric field. Fig. 2(c) shows the linewidth of bound exciton emission as a function of voltage. We extract the linewidth by calculating the full width at half maximum of the Voigt fit. For voltages below 35 V, the linewidth remains constant and is limited by the spectrometer resolution. This range, in good agreement with Fig. 2b, is the regime in which the device operates at optimal capacity. Once the applied voltage exceeds 35 V, we observe linewidth start to broaden, which we attribute to increased sensitivity of the emitter to background charge fluctuations [DeSantis2021Stark] and possible heating from the electrodes.
While high fields eventually induce broadening, we observe an intermediate regime where the field significantly reduces the bound exciton linewidth by actively suppressing spectral diffusion. We investigate this stabilization mechanism in the following section
IV Manipulation of Charge Environment
To isolate the effect of the electric field on the linewidth, we need to change the excitation method. Linewidth is strongly determined by local charge environment, and above-band pumping generates a significant background of free carriers, perturbing the local charge landscape. Therefore, photoluminescence spectroscopy cannot show any subtle effects of electric-field-induced charge stabilization around the emitter. Although directly driving the bound-exciton resonance is possible through the observed polarization selectivity (see Supplemental Material section II), cross-polarization is limited due to unwanted reflection from metal surfaces.
As an alternative approach, we perform photoluminescence-excitation spectroscopy by resonantly pumping the free exciton and measuring the bound exciton emission. A narrow-linewidth laser ($\approx 90\mathrm{kHz}$) is tuned to selectively pump the free exciton on resonance, which induces the bound exciton emission through a cascaded process. Before each measurement, we compensate for the Stark shift of the free exciton by tuning our laser back on the free-exciton resonance (see Supplemental Material section III.B)
Fig. 3(a) shows the photoluminescence-excitation spectra for different values of voltage, all centered at the same wavelength. In the absence of an electric field, the signal initially exhibits about three times broader linewidth than the spectrometer limited resolution photoluminescence signal. In the presence of the field, the linewidth undergoes a clear reduction as the electric field increases. We repeat this measurement over a broader range of applied voltages while keeping the resonant power fixed ($50\text{\,}\mathrm{n}\mathrm{W}$). Fig. 3(b) shows the extracted linewidth as a function of applied biases. We extract the linewidth values from a Voigt model fit to the spectrum at each bias and then normalize them to the zero-field intensity value. The trend shows an initial narrowing dynamic followed by a monotonic broadening effect. This behavior reveals an optimal field around $35\text{\,}\mathrm{V}$, at which the emission linewidth reaches a minimum at $50\%$ of the initial linewidth. Beyond this point, the observed broadening is consistent with previous photoluminescence observations, showing sensitivity to background charge fluctuations. The initial broadening and its subsequent suppression with applied field can be attributed to trap states [karczewski1994deep, hernandez1996investigation] and suppression of the resulting charge fluctuations in the emitter’s local environment [Anderson2019Electrical, Somaschi2016NearOptimal, houel2012probing]. The dashed curve shows the best fit based on a statistical model, elaborated upon later in this section.
As further evidence on the suppression of trap states, we perform photoluminescence excitation spectroscopy while adding a weak above-band pump in the absence of an electric field. The power level used in this experiment is approximately two orders of magnitude lower than the saturation level (see Supplemental Material section V), resulting in negligible direct emission from above-band. Fig. 3(c) shows the photoluminescence-excitation signal for different values of the above-band power, all centered at the same wavelength. The linewidth exhibits clear narrowing as the above-band power increases. We repeat the measurement across a broader span of above-band powers while keeping the resonant power fixed. Fig. 3(d) shows the extracted central wavelength (black dots) and normalized linewidth (blue dots) as a function of weak above-band power. The emitter is initially broad and undergoes a clear reduction in linewidth until it reaches the spectrometer’s limited resolution, almost matching the photoluminescence signal and reaching a final value around $40\%$ of initial linewidth. Concurrently, the central wavelength shifts toward the blue in response to increasing power, reaching a saturation plateau. We extract the central wavelength and linewidth through a Voigt fit to the spectra at each above-band power and normalize linewidths to the zero-above-band power value. We attribute both observations to the generation of excess carriers under above-band illuminations that fill nearby trap states. This effect results in suppression of the built-in DC Stark effect and electric-field fluctuations, manifesting as reduced spectral diffusion and linewidth narrowing [Majumdar2011-yp, Gazzano2018-ki, Chen2016-pj, PhysRevB.93.195316]. The dashed curves show the extracted best fit for relative linewidth and central wavelength based on a statistical model described in the subsequent discussion. We note that in previous studies [Karasahin2022, Jiang2024], a weak above-band pump was necessary in order to observe any photoluminescence excitation signal from the donor-bound exciton. The donors were initially ionized, and the above-band light was required to supply the electron ground state. Observing the direct photoluminescence excitation signal here, without such pumping, suggests that the surface passivation and oxide capping are effective in partially stabilizing the donors.
[…]
$\mu_{\delta\omega}=$ $\beta\Big(E_{0}^{2}+p\,S_{2}\Big)$ (1)
$\sigma_{\delta\omega}^{2}=$ $\beta^{2}\left[p(1-p)\big(S_{4}+2E_{0}^{2}S_{2}\big)\;+\;p^{2}(1-p^{2})\big(S_{2}^{2}-S_{4}\big)\right]$ (2)
where $p=k^{+}/(k^{-}+k^{+})$ is the probability of trap occupation, $f_{i}$ is the field of $i$th trap in the donor site appeared in geometric moments of the trap fields $S_{n}=\sum_{i=1}^{N}f_{i}^{n}$, $N$ is the number of contributing traps, $E_{0}$ is the electric field bias, and $\beta$ is the polarizability of the emitter.
Our model qualitatively explains the stabilization dynamics of trap states by framing them as a competition between carrier capture and release rates. Based on this model, charge fluctuation is diminished when either rate dominates, pushing the occupation probability ($p$) toward the stable limits of 0 or 1. Under electrical suppression, band bending increases the tunneling rate [ganichev2000distinction], driving the occupation probability $p$ toward 0, illustrated in Fig. 3(e). At higher fields, the bias-noise coupling term $2E_{0}^{2}S_{2}$ begins to dominate, leading to monotonic spectral broadening and thus defining a sweet spot for linewidth minimization, reflecting the behavior in Fig. 3(b). In the case of optical suppression, above-band pumping increases the capture rate via photo-generated carriers [PhysRev.87.835], pushing $p$ toward 1, which is illustrated in Fig. 3(f). Increased occupation leads to a blue-shifted central wavelength and a narrower emission linewidth, as shown in Fig. 3(d). Additional detail about the extracted fit parameters, derivation and Monte Carlo simulation based on this model is provided in Supplemental Material section VI.
V Conclusion and outlook
We show that laterally applied electric fields in ZnSe quantum wells enable local tuning of single impurity–bound exciton emission energies, compensating for spectral inhomogeneity arising from strain and charge environment imperfections. Beyond spectral control, these fields can suppress trap-induced charge fluctuations, resulting in observable linewidth narrowing for the donor-bound exciton transition. We also verify this through weak above-band optical pumping and a statistical model, qualitatively describing the observed phenomena. Based on current results, the precise tuning and reduction of spectral wandering may lead to higher visibility in Hong–Ou–Mandel (HOM) from two separate devices, addressing previous challenges. Additionally, it can potentially lead to enhanced spin coherence times. Looking ahead, embedding the quantum well in a Schottky diode structure would introduce an out-of-plane field that directly addresses the vertical confinement axis, offering stronger control and the potential to restore transform-limited linewidths. Together, these observations establish electric fields as a versatile approach for both finely tuning emission energies and stabilizing the charge environment in impurity-bound excitonic emitters.
Acknowledgments
The authors acknowledge Shahriar Aghaei, Kai-Mei Fu, Douglas L Irving and Kelsey J Mirrielees for valuable discussions. The Waks group would like to acknowledge support from the AFOSR (grant #FA95502010250) and the Maryland-ARL quantum partnership (grant #W911NF1920181). The Pawlis group would like to acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - Cluster of Excellence Matter and Light for Quantum Computing (ML4Q) EXC 2004/1-390534769. The funder played no role in study design, data collection, analysis and interpretation of data, or the writing of this manuscript.
DATA AND MATERIALS AVAILABILITY
All of the data that support the findings of this study are reported in the main text and Supplementary Materials. Source data are available from the corresponding authors on reasonable request.
Competing interests
The authors declare that they have no competing interests.
Author contributions
A.A. and E.W. conceived the experiment. C.F., N.D., and Y.K. fabricated the quantum well and performed initial characterization. A.A. patterned the electrodes and made the connections. A.A. performed the experiment. Y.J. and A.B. supported setting up the experiment. A.A. analyzed the data. A.A., J.B., and E.W. prepared the manuscript. All authors discussed the results and confirmed the manuscript. E.W. and A.P. supervised the experiment.